# Generative Predictive Control: Flow Matching Policies for Dynamic and Difficult-to-Demonstrate Tasks

### v1

@@ VI Domain Randomization Strategies
Fortunately, the availability of massively parallel simulators and the structure of the SPC/GPC paradigm enables a range of new DR possibilities. In particular, we can modify the SPC rollouts (4) by simulating each action sequence $U^{(i)}$ in several domains with randomized parameters (e.g., friction coefficients, body masses, etc.). This results in cost values indexed by both sample $i$ and domain $d$, e.g.,
@@
We then aggregate this cost data across domains before performing the standard SPC update (5) on the net costs.
The simplest choice would be take the average cost over all domains,
@@
But the SPC/GPC framework allows for other possibilities as well. We can, for instance, take a more conservative approach and use the worst-case cost
@@
We can even use more sophisticated risk metrics like conditional value-at-risk (CVaR) [39, 40], which takes the expected cost in the $(1-\beta)$ tail of the distribution:
$J^{(i)}=\inf_{z\in\mathbb{R}}\mathbb{E}_{d}\left[z+\frac{\max(J^{(i,d)}-z,0)}{1-\beta}\right].$ (39)
The parameter $\beta\in[0,1]$ determines the degree of risk sensitivity.
These and other risk strategies for online domain randomization are implemented in hydrax [21], making them readily available for both GPC training and online deployment in a bootstrapped SPC controller. We illustrate the impact of different domain randomization strategies in Section VII-D below.
VII Simulation Experiments
Figure 3: Systems used to evaluate GPC performance, from left to right: inverted pendulum, cart-pole, double cart-pole, push-T, planar biped, luffing crane, humanoid.
In this section, we present the results from simulation experiments on each of the systems shown in Fig. 3. We aim to answer the following questions:
  1. 1.
    Can GPC perform tasks that require multi-modal reasoning, as well highly dynamic tasks that require high-frequency feedback (Sec. VII-B)?
  2. 2.
    Can GPC continually improve its performance over multiple iterations (Sec. VII-C)?
  3. 3.
    How do different domain randomization strategies impact performance with model error (Sec. VII-D)?
  4. 4.
    What are the scalability limits of this approach (Sec. VII-E)?
We find that GPC is effective for control systems with fast dynamics at high feedback rates, enjoys the training stability characteristic of supervised learning methods, and enables risk-aware control, but struggles to scale to our largest and most difficult example (humanoid standup). We provide further discussion of these scalability limitations, and how they might be overcome in the future, in Section VIII.
VII-A Example Systems
We test GPC on seven example systems of varying scale and difficulty. Each of these systems are introduced briefly below, with further details available in the open-source implementation [34].
Inverted pendulum: This simple one-dimensional system requires swinging a pendulum to the upright position and balancing it there. Torque limits prevent the pendulum from swinging directly upright: the policy must gradually pump energy into the system.
Cart-pole: An unactuated pendulum is mounted on an actuated cart. Control actions are torques applied to the cart, and the task is to balance the pendulum upright. While this is a relatively simple nonlinear system, obtaining successful human demonstrations would be difficult.
Double cart-pole: In this extension of the cart-pole example, an unactuated double pendulum is mounted on the cart. The fast and chaotic double pendulum dynamics make this task particularly challenging.
Push-T: A robotic finger pushes a T-shaped block to a target position and orientation on a table. This task has been solved with behavior cloning [1], and is a standard example of a task that requires multi-modal reasoning.
Planar biped: A robot walker, constrained to the sagittal plane, is tasked with moving forward at 1.5 m/s. The high dimensionality of this system would make teleoperation for behavior cloning difficult. Successful locomotion also requires fast replanning.
Luffing crane: A swinging payload is attached via a rope to a luffing crane. Control actions are target boom angles and rope length. This underactuated system provides a particularly useful testbed for investigating the impact of modeling errors and domain randomization strategies.
Humanoid standup: A Unitree G1 humanoid model is tasked with reaching a standing configuration. Initial conditions are random joint angles, joint velocities, and base orientation, so that the robot begins sprawled on the ground.
A brief summary of each of these examples is shown in Table I. For the smaller examples we use a simple Multi-Layer Perceptron (MLP) architecture for the flow network $v_{\theta}$, while for the larger examples we use a Convolutional Neural Network (CNN) with FiLM conditioning [41], as recommended in [1]. All experiments were performed on a desktop computer with an Nvidia RTX 4070 (12 GB) GPU.
Name Pendulum Cart-Pole Double Cart-Pole Push-T Walker Crane Humanoid
@@ VII-A Example Systems
Planning horizon (sec) 0.5 1.0 0.8 0.5 0.6 0.8 0.9
Network architecture MLP MLP MLP CNN CNN CNN CNN
Train time (hr:min:sec) 0:00:30 0:01:20 0:17:27 0:17:22 0:06:25 0:09:41 2:46:55
Inference time (ms) 1.0 1.3 1.3 3.8 8.1 2.4 4.0
GPC-seeded SPC is effective
GPC alone is effective
TABLE I: Summary of each of the example systems. Full details and hyperparameter configurations are available at [34]
VII-B Policy Performance
Footage of closed-loop GPC performance on each of the examples is shown in the accompanying video. In all the examples, the GPC policy was effective in improving closed-loop performance when used in conjunction with SPC, as described in Sec. V-B. Applying the GPC policy directly (Sec. V-A) was effective in all cases except the humanoid.
GPC can handle multi-modal action distributions. This is evidenced by effectiveness on the Push-T example, which requires multi-modal reasoning to reach around the block [1]. Interestingly, GPC training takes under 20 minutes, while training a similar Push-T policy with diffusion-based behavior cloning (and full state observations) takes around an hour.
Figure 4: Closed-loop performance for the double cart-pole with and without warm-starts. The warm-started GPC controller (right) produces smooth control actions and is able to successfully balance the double pendulum upright. Without warm-starts (left), control actions are dominated by noisy jittering between modes, and the system does not balance.
More importantly, GPC can control systems with fast dynamics at high control rates. The double cart-pole example provides a particularly clear example of this, as well as the importance of warm-starts (Sec. V-A). Fig. 4 illustrates performance with and without warm-starts. Without warm-starts (left, $\alpha=0$), the control actions are dominated by significant noise (top plot) and the system is unable to swing upright (bottom plot). With warm-starts (right, $\alpha=1$), we obtain smoother control actions and successfully balance around the upright position. The controller is still able to respond rapidly to the chaotic system dynamics, as evidenced by the rapid changes between 1 and 2 seconds.
VII-C Training Stability
Training curves for each of the examples are shown in Fig. 5. In addition to the average loss and average cost at each iteration, we plot the percent of states in which the flow-matching policy generates the best action sequence. Recall that some samples come from the policy, while others are from the SPC normal proposal distribution (Algorithm 1, lines  1-1). A higher percentage of best samples coming from the policy is a useful indicator of policy performance, and tends to increase monotonically (modulo noise from randomized initial conditions) throughout the training process.
@@ VII-C Training Stability
Figure 5: Training curves showing the average loss $\mathcal{L}_{GPC}$, percent of states in which the flow-matching policy generated the best action sequence, and the average cost of the best rollout. GPC is able to leverage the training stability of supervised learning while avoiding the need for demonstrations.
@@
Interestingly, this percentage does not reach 100%, even after many iterations. We suspect that this is due to the fact that a good sample from the policy also improves SPC performance at the following time step, since this will shift the mean of the proposal distribution, $\bar{U}_{k}$, close to optimality.
While we leave a systematic hyperparameter sensitivity study for future work, our empirical observation is that GPC benefits from the training stability characteristic of supervised learning methods. This is in contrast to reinforcement learning methods, which can exhibit high sensitivity to reward tuning, implementation details, and even the random seed used for training [9, 8].
VII-D Domain Randomization
GPC offers an opportunity to train with risk-sensitive domain randomization schemes, as outlined in Sec. VI. We use the luffing crane example to explore the impact of different strategies on the performance of a trained policy.
In particular, we train three GPC policies: one with no domain randomization, one with average-cost domain randomization (37), and one with more conservative CVaR ($\beta=0.25$) domain randomization (39). For the domain randomized models, we used 8 randomized domains. Each domain uses a slightly different model with modified joint damping, payload mass, payload inertia, and actuator gains.
After training, we apply each policy directly with warm-starts (Sec. V-A). To evaluate closed-loop performance, we generate 50 random target locations that the payload must visit. The target is moved to the next location once the payload is moved within 15 cm or after 10 seconds, whichever comes first. We report the total time to visit all 50 targets in Table II: lower times indicate better performance. The same target locations are used for each policy.
No DR Avg. CVaR(0.25)
No model error 106 103 133
With model error 165 184 139
TABLE II: Time (in seconds) for the crane to visit a sequence of 50 randomly generated payload target positions, under policies trained with various domain randomization schemes.
We first test the policies without any model error in the simulator. In this case both the non-randomized policy and average-case domain randomization perform significantly better than the more conservative CVaR policy. When we add a small amount of model error to the simulator (lower joint damping, heavier payload mass), all methods perform worse. But the more conservative CVaR policy degrades the least, and significantly outperforms the other policies.
VII-E Scalability
In choosing our set of example systems, which range from 1 to 29 degrees-of-freedom, we aim to push the limits of the scalability of our method. This scalability hits a bottleneck with the largest humanoid example: the trained policy is unable to reliably drive the robot to stand up, though using the policy to seed SPC is effective (as shown in the accompanying video).
It is likely that further cost and hyperparameter tuning, as well as the addition of more computation, would result in better performance for the humanoid example. However, state-of-the-art reinforcement learning methods are able to train similar behaviors [42, 43], suggesting that further algorithmic improvements in the GPC framework are necessary to achieve reliable scalability to such systems.
VIII Limitations and Future Work
Limited effectiveness on our largest (humanoid standup) example is the most severe limitation of our proposed method, as discussed above. We believe that value function learning will be a critical component in overcoming these scalability limitations. In addition to being a key element of state-of-the-art reinforcement learning methods like PPO, value learning would allow us to reduce the planning horizon $T$ in problem 1. Reducing the planning horizon reduces the dimensionality of the sampling space, making planning easier while maintaining long-horizon reasoning. Methods that leverage connections between the gradient of the value function and the flow field $v_{\theta}$ are of particular interest.
Similarly, the sample complexity of the basic GPC framework introduced here is poor relative to state-of-the-art RL algorithms. To generate a single training data point, we run $N$ short simulations—and even more under a risk-based domain randomization strategy. While advances in computing speeds may mean that the benefits of a supervised learning framework eventually outweigh the computational costs, methods of more fully using all of the data from SPC rollouts could significantly improve the GPC framework.
Other performance improvements could come from more benign algorithmic details. For instance, our SPC implementation [21] represents action sequences with a simple zero-order-hold spline parameterization. Higher-order splines [18] or alternative parameterizations could be more effective. Better choices of action space, such as using task-space/end-effector coordinates rather than joint coordinates, could also be useful. Actuation limits are a critical component of many robotics tasks, but our implementation does not handle these in any particularly special way. Leveraging recent advances in constrained generative modeling [44, 45, 46] to do so is another potentially fruitful area for future work.
Future work will also focus on hardware experiments. Hardware experiments are a critical component of evaluating any robotic control framework: this is particularly the case when it comes to domain randomization strategies. This will also provide an important platform for exploring policies conditioned on various observations, ranging from raw sensor data to images to foundation model embeddings like DINO [47]. Integrated simulation and rendering in the recently-released MuJoCo playground [48] could provide a useful software platform for exploring image-conditioned policies.
Finally, the GPC framework offers a means for including dynamic and difficult-to-demonstrate tasks in a generalist policy [2] or large behavior model [10] framework that combines data from multiple tasks in a single policy. Exploring this possibility remains an important topic for future work.
IX Conclusion
We introduced Generative Predictive Control, a framework for learning flow matching policies on dynamic tasks that are easy to simulate but difficult to demonstrate. GPC leverages tight connections between generative modeling and sampling-based predictive control to generate training data for supervised learning without expert demonstrations. We showed how warm-started GPC policies enable real-time control at over 500 Hz, ensuring temporal consistency via warm-starts. Future work will focus on validating the GPC framework on hardware, incorporating value function learning, and training generalist multi-task policies.

### v2

@@ Appendix D Appendix: Domain Randomization Strategies
The availability of massively parallel simulators and the structure of the SPC/GPC paradigm enables a range of new DR possibilities. In particular, we can modify the SPC rollouts (4) by simulating each action sequence $U^{(i)}$ in several domains with randomized parameters (e.g., friction coefficients, body masses, etc.). This results in cost values indexed by both sample $i$ and domain $d$, e.g.,
@@
We then aggregate this cost data across domains before performing the standard SPC update (5).
The simplest choice for aggregation would be take the average cost over all domains,
@@
But the SPC/GPC framework allows for other possibilities as well. We can, for instance, take a more conservative approach and use the worst-case cost,
@@
Or we can use more sophisticated risk metrics like conditional value-at-risk (CVaR) [37, 38], which takes the expected cost in the $(1-\beta)$ tail of the distribution:
$J^{(i)}=\inf_{z\in\mathbb{R}}\mathbb{E}_{d}\left[z+\frac{\max(J^{(i,d)}-z,0)}{1-\beta}\right],$ (35)
where $\beta\in[0,1]$ determines the degree of risk sensitivity.
These and other risk strategies for online domain randomization are implemented in hydrax [24], making them readily available for both GPC training and online deployment in a bootstrapped SPC controller. Section 6.3 illustrates the impact of different domain randomization strategies on the luffing crane example.
Appendix E Appendix: Example System Details
Details regarding each of the example systems from Section 6 can be found below, with further details in Table 2. We represent the action sequence $U$ with a zero-order-hold spline over a fixed number of knot points. Using few knot points is critical for good SPC performance, particularly on the higher-DoF examples.
Pendulum Cart-Pole
Double
Cart-Pole
Push-T Walker Crane Humanoid
@@ Appendix E Appendix: Example System Details
Observation Size 3 5 8 5 18 23 56
Planning horizon (s) 0.5 1.0 0.8 0.5 0.6 0.8 0.9
Planning knots 5 10 10 5 4 2 3
Table 2: Example task specification parameters.
Inverted pendulum: This simple one-dimensional system requires swinging a pendulum to the upright position and balancing it there. Torque limits prevent the pendulum from swinging directly upright: the policy must gradually pump energy into the system. The observation $y=h(x)$ consists of sine and cosine of the angle, along with angular velocity.
Cart-pole: An unactuated pendulum is mounted on an actuated cart. Control actions are torques applied to the cart, and the task is to balance the pendulum upright. While this is a relatively simple nonlinear system, obtaining successful human demonstrations would be difficult. The observations are sine and cosine of the pendulum angle, position of the cart, and linear and angular velocities.
Double cart-pole: In this extension of the cart-pole example, an unactuated double pendulum is mounted on the cart. The fast and chaotic double pendulum dynamics make this task particularly challenging. As for the cart-pole, observations include sine and cosine of pendulum angles, cart position, and linear and angular velocities.
Push-T: A robotic finger pushes a T-shaped block to a goal position and orientation on a table. This task has been solved with behavior cloning [1], and is a standard example of a task that requires multi-modal reasoning. The observations include pusher position, block position, and block orientation.
Planar biped: A robot walker, constrained to the sagittal plane, is tasked with moving forward at 1.5 m/s. The high dimensionality of this system would make teleoperation for behavior cloning difficult. Successful locomotion also requires fast replanning. The observation is the full system state (positions and velocities), excluding the horizontal position.
Luffing crane: A swinging payload is attached via a rope to a luffing crane. Control actions are target boom angles and rope length. This underactuated system provides a particularly useful testbed for investigating the impact of modeling errors and domain randomization strategies. The observation includes crane joint angles and the position and velocity of the payload relative to the target.
Humanoid standup: A Unitree G1 humanoid model is tasked with reaching a standing configuration. Initial conditions are random joint angles, joint velocities, and base orientation, so that the robot begins sprawled on the ground. The observation includes all joint angles and joint velocities, floating base velocity, floating base position, and floating base orientation relative to the upright.
Appendix F Appendix: Training Details
Hyperparameters for policy training and inference are shown in Table 3. When evaluating a trained SPC or GPC+ policy, we used a total of 128 simulation rollouts. For SPC, all of these are drawn from a Gaussian proposal distribution. For GPC+, 64 are from the Gaussian distribution and 64 are from the flow matching policy.
Pendulum Cart-Pole
Double
Cart-Pole
Push-T Walker Crane Humanoid
Architecture MLP MLP MLP CNN CNN CNN CNN
Iterations 10 10 50 20 20 10 20
Train (min:sec) 00:30 01:20 17:10 17:30 9:36 6:15 113:18
Envs. ($N_{E}$) 128 128 256 128 128 512 128
SPC samp. ($N_{S}$) 8 8 16 128 16 4 32
Policy samp. ($N_{P}$) 2 2 16 32 16 2 32
Episode len. (sec) 4.0 2.0 4.0 4.0 5.0 5.0 8.0
Num. Params. 1541 5898 20362 19990 49768 48680 368567
Batch size 128 128 128 128 128 1024 128
Learning rate $10^{-3}$ $10^{-3}$ $10^{-3}$ $10^{-3}$ $10^{-3}$ $10^{-3}$ $10^{-3}$
Epochs 10 100 10 10 10 20 10
Inference (ms) 1.0 1.3 1.3 3.8 8.1 2.4 4.0
ODE step size $\delta t$ 0.1 0.1 0.1 0.1 0.1 0.1 0.1
Ctrl. Freq. (Hz) 50 50 50 50 50 30 50
Table 3: Summary of training and inference hyperparameters.
The hyperparameters in Table 3 are not highly optimized: our primary objective is merely to confirm the ability of GPC to train usable generative policies. We leave a systematic study of the relative impact of various hyperparameters for future work.
@@ Appendix F Appendix: Training Details
Figure 5: Training curves showing the average cost $J$, percent of states in which the flow-matching policy generated the best action sequence, and the loss $\mathcal{L}_{GPC}$ from three random seeds. GPC is able to leverage the training stability of supervised learning while avoiding the need for demonstrations.
Training curves for each of the examples are shown in Fig. 5. In addition to the average loss and average cost at each iteration, we plot the percent of states in which the flow-matching policy generates the lowest-cost action sequence. Recall that some samples come from the policy, while others come from the SPC normal proposal distribution (Algorithm 1, lines  1-1). A higher percentage of best samples coming from the policy is a useful indicator of policy performance, and tends to increase (modulo noise from randomized initial conditions) throughout the training process.
@@
Interestingly, this percentage does not reach 100%, even after many iterations. This is due to the fact that a good sample from the policy also improves SPC performance at the following time step, since a good policy sample at step $k-1$ shifts the mean of the proposal distribution at step $k$ close to optimality.